Circular No. 8096

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

IGR J16320–4751 J. Rodriguez, Commissariat à l’Energie Atomique (CEA), Service

d’Astrophysique (SA), Saclay; J. A. Tomsick, University of California, San Diego (UCSD); L. Foschini, Instituto di Astrofisica Spaziale e Fisica Cosmica, Consiglio Nazionale delle Ricerche, Bologna; R. Walter, Integral Science Data Center, Versoix; A. Goldwurm, CEA/SA, report that an XMM-Newton target-of-opportunity observation on Mar. 4.85 UT shows a bright source at α = 16^{h}32^{m}01^{s}.9, δ = −47^{o}52′29′′ (equinox 2000.0; uncertainty ∼ 4′′), with both EPIC cameras (pn and MOS) during an effective exposure of ∼ 4500 s. The source location is consistent with that for IGR J16320– 4752 and AX J1631.9–4752 reported from earlier INTEGRAL, ASCA, and Beppo-SAX observations (IAUC 8076, 8077; Sugizaki et al. 2001, Ap.J. Suppl. 134, 77). A fit to an EPIC-pn spectrum with an absorbed power law gives a photon index of 1.6_{−0.1}^{+0.2} and a hydrogen column density of 21.3_{−1.4}^{+3.7} × 10^{22} cm^{−2} (1σ uncertainties). The EPIC-pn source count rate varies between ∼ 0.1 and 0.6 count/s during the observation. The difference in the power-law index, compared to that obtained with ASCA and Beppo-SAX, may suggest that the source undergoes spectral transitions similar to those of x-ray binaries. Longer x-ray observations and multiwavelength follow-ups are strongly encouraged.

Tomsick, Rodriguez, Goldwurm, Walter, and Foschini, also write: “Infrared images from the 2MASS database show that two sources are present within the XMM-Newton error circle for IGR J16320–4751 = AX J1631.9– 4752 (IAUC 8076, 8077). The position of the source (with magnitudes K_{s} = 10.9 ± 0.1, H = 12.9 ± 0.2, J > 16) that is closest to the center of the x-ray error circle is α = 16^{h}32^{m}01^{s}.75, δ = −47^{o}52′28′′.9 (equinox 2000.0; uncertainty 1′′). The second source (K_{s} = 10.8, H = 11.2, J = 12.1) is on the southeast edge of the x-ray error circle at position end figures 02^{s}.18, 32′′.1. For the first source, the colors suggest a relatively high level of extinction, which is compatible with the high neutral hydrogen column density inferred from x-ray observations of this source. This, along with the better positional agreement, makes the first source a more likely candidate. Infrared observations are encouraged to check the identification.”

COMET C/2001 RX_{14} (LINEAR) Visual m_{1} estimates: Feb. 23.90 UT, 10.3 (K. Hornoch, Lelekovice,

Czech Rep., 0.13-m reflector); Mar. 4.01, 11.0 (J. Carvajal, Madrid, Spain, 0.21-m refl.); 13.17, 11.1 (A. Diepvens, Balen, Belgium, 0.15-m refractor).

2003 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green